RC meeting and Run Plan, Tuesday October 26th

From SBS wiki
Jump to: navigation, search

Meetings Page

RC Meeting Information

RC Daily Meetings are at 3:15pm, remote connection link: https://bluejeans.com/256205934

Accelerator Status

  1. DC Power still in tunnel
  2. Pass Change not started
  3. 3.5 uA limit to Hall A due to BLM trips. Higher currents require negotiations with Hall C

Hall A Status and News

  1. In Restricted Access
  2. Move to SBS-5 in progress

Current Run Plan

  1. Pass-4 Setup
  2. SBS-5

Subsystem Updates

  1. HCAL
  2. BBCal
  3. Timing Hodoscope
  4. Grinch
  5. GEMS
  6. DAQ
  7. Replay Analysis

Reminders to shift crew

  1. Check GEM gas mix every 2 hours
  2. HALOG GEM efficiency plots e.g. (https://logbooks.jlab.org/entry/3930258)
  3. Fill Out Run list
  4. New Quick Beam checkout procedure (Target at home, Harps SCAN, check BPM strip chart) No C Hole after initial checkout completed at each configuration
  5. Ramp BB magnet to 0 before "power off". Here is a nice entry by Jack here

Current Issues

  1. Can't identify pulsar events in HCal (Ole, Alex)
  2. Need script/method to determine # good events
  3. Need script to get accumulated charge
  4. Waiting for response from OPS about when "long" ramping procedure required on SBS quick ramp vs [ https://logbooks.jlab.org/entry/3931987 long procedure]
  5. Rates in Hodoscope at ("300 kHz/1 uA on LD2 signals, 10 mV threshold, 40 mV signal on "FADC", "good" signal based on GeM tracks" D.Hamilton) Would like to do a HV scan on timing hodoscope (Ralph)

Special Requests

Useful Links